### v3

@@ Abstract
Most existing methods rely on the similarity correlation matrix to characterize the repetitiveness of actions, but their scalability is hindered due to the quadratic computational complexity. In this work, we introduce a novel approach that employs an action query representation to localize repeated action cycles with linear computational complexity. Based on this representation, we develop two key components to tackle the essential challenges of temporal repetition counting. Firstly, to facilitate open-set action counting, we propose the dynamic update scheme on action queries. Unlike static action queries, this approach dynamically embeds video features into action queries, offering a more flexible and generalizable representation. Secondly, to distinguish between actions of interest and background noise actions, we incorporate inter-query contrastive learning to regularize the video representations corresponding to different action queries. As a result, our method significantly outperforms previous works, particularly in terms of long video sequences, unseen actions, and actions at various speeds. On the challenging RepCountA benchmark, we outperform the state-of-the-art method TransRAC by 26.5% in OBO accuracy, with a 22.7% mean error decrease and 94.1% computational burden reduction. Code will be publicly available.
@@ 1 Introduction
Temporal periodicity is a ubiquitous phenomenon in the natural world. Temporal Repetition Counting (TRC) aims to accurately measure the number of repetitive action cycles within a given video and holds significant potential for applications such as fitness monitoring [11] and motion generation [49].
@@
Pioneer methods  [17, 1, 7, 39, 29, 38, 6] represent time-series video data as one-dimensional signals and employ spectral analysis techniques such as the Fourier transform. While suitable for short videos with fixed periodic cycle lengths, these methods struggle to handle real-world scenarios with varying cycle lengths and sudden interruptions. Recent studies shift to deep learning-based methods [18, 9, 14, 47, 19] and show promising performances. Notably, most of these methods, such as RepNet [9] and TransRAC [14], utilize a temporal similarity correlation matrix to depict repetitiveness, as illustrated in Fig. 1 (a). Nevertheless, the computational complexity of this representation grows quadratically with the number of input frames $T$, highlighting a significant gap in scalability that hinders their application to real-world scenarios of varying action periods and dynamics.
Recent progress in action detection [34, 24] introduces an efficient representation of action periods by associating each action instance with an action query, similar to DETR [3]. Inspired by this, we propose to formulate the TRC problem as a set prediction task where the goal is to detect every action cycle by representing it as an action query, as illustrated in Fig. 1 (b). Based on this query representation, we use a Transformer encoder-decoder network [50] to detect repetitive action instances and their temporal positions, defined by their midpoints and durations. This novel formulation reduces the complexity from quadratic to linear  [^1] In implementation, we employ deformable attention modules as proposed in DeformableDETR [50]. and enables counting long videos with varying action periods. However, directly applying the action detection approaches [24, 45, 34] to the TRC problem proves inadequate (Tab. 1) in addressing two distinctive challenges unique to TRC. These challenges underscore the complexity of TRC, highlighting why TRC is not merely another action detection task but requires a nuanced approach that considers the unique nature of repetitive actions. We highlight the two inherent differences between TRC and the classical action detection task:
@@
As a result, approaching TRC as a simple action detection task results in inferior performance as shown in Sec. 4.3. In contrast, we propose two novel strategies to address these challenges and redefine the framework for TRC. In response to the first challenge, we propose Dynamic Action Query (DAQ), which adaptively updates the action query using content features extracted from the video (Fig. 2). This mechanism allows the decoder to attend to the repetitive actions based on the input video contents in a dynamic and contextually aware manner, without the need for manual definition. To tackle the second challenge, we define two action classes: “repetitive actions”, which are dynamically defined by DAQ, and the “others ($\emptyset$)” class. We further propose Inter-query Contrastive Learning (ICL) which ensures that primary repetitive action cycles are grouped together in the learned representation space while being separated from other distractors, such as background noise. The integration of two core components (DAQ and ICL) ensures that the action instances are identified adaptively based on the video content and their contextual similarity. In other words, our queries are designed to localize the contextually similar action instances, which aligns exactly with the definition of repetition counting. Extensive experiments validate the effectiveness of the two proposed designs.
@@
    We provide a novel perspective to tackle the TRC problem using a simple yet effective representation for action cycles. Our approach reduces the computational complexity from quadratic to linear and is robust to varying action periods and video lengths.
@@
    By addressing the scalability issue and unique challenges of TRC, our method on two challenging benchmarks notably surpasses state-of-the-art (SOTA) methods in terms of both accuracy and efficiency. Notably, our method strikes an effective balance in handling various action periods and video lengths, offering a significant leap forward in the practical application of TRC technologies.
@@ 2.1 Temporal Repetition Counting
Traditional methods [17, 1, 7, 39, 29, 38, 6] frequently employ spectral or frequency domain techniques for the analysis of repetitive sequences, thereby preserving the underlying repetitive motion structures. While these conventional approaches are capable of effectively handling simple motion sequences or those characterized by fixed periodicity, they prove inadequate when confronted with non-stationary motion sequences encountered in real-world scenarios. In contrast, deep-learning-based approaches [18, 9, 14, 47, 19] have demonstrated remarkable performance improvements. Notably, RepNet [9] and TransRAC [14] leverage temporal similarity matrices of actions to construct models for counting temporal repetitions. However, these similarity-matrix-based methods are not scalable for long videos due to their quadratic computational complexity. Another research line involves predicting the start and end points of each cycle [47] from coarse to fine. Nevertheless, its practicality is hindered by the requirement for over 30 forward passes to count iteratively from a single video. In this paper, we introduce an effective action cycle representation by leveraging a Transformer encoder-decoder, which reduces the computational complexity from quadratic to linear and demonstrates superior performance in handling both fast and slow actions.
@@ 2.2 Temporal Action Detection
The field of temporal action detection [31, 48, 5, 45, 22] is typically classified into two categories: anchor-based methods, and anchor-free methods. Anchor-based methods [44, 20, 30] generate multiple anchors, subsequently classifying these anchors to determine the action boundaries. Anchor-free methods  [2, 35, 43, 21] predict action instances by directly regressing the boundary and the center point of an action instance. With the rapid development of Transformer technology, DETR [3] is introduced for object detection task [50, 28, 23, 46] and gains increasing popularity with promising performance. This paradigm promotes the study in many fields such as the action detection tasks [24, 37, 40, 42]. These methods establish a direct connection between action queries and the predicted action instances, enabling them to accurately predict the temporal boundaries of actions. Inspired by these promising results, we explore the possibility of utilizing a novel action query to represent the action cycle in TRC task. In contrast to existing action detection methods, our approach allows the model to capture the inherent repetitive content of an action cycle without relying on predefined class labels and effectively addresses confounding factors such as non-repetitive video backgrounds. This makes our approach well-suited for tackling the challenges of the TRC problem.
@@ 3.1 Preliminary
DETR [3] is a pioneering object detection framework that builds upon the Transformer encoder-decoder architecture [40]. The overall DETR architecture [3] consists of three main components: a backbone to extract image features, an encoder-decoder Transformer, and the detection heads, i.e. feed-forward network (FFN) that makes the final detection prediction. The main features of DETR are the conjunction of a bipartite matching loss and transformers with (non-autoregressive) parallel decoding. The bipartite matching loss is a set-based Hungarian loss during training, which uniquely assigns a prediction to a GT object, and is invariant to a permutation of predicted objects. This design enables DETR to perform parallel processing and predict all objects simultaneously. We briefly review the workflow as follows.
Given the input image feature maps extracted by a CNN backbone, e.g. ResNet [13], DETR exploits a standard Transformer encoder-decoder network to transform the feature maps to be features of a set of object queries. An FFN and a linear projection are added on top of the object query features as the detection heads. The FFN acts as the regression branch that predicts the bounding box coordinates, i.e. box center coordinates, box height and width. The linear projection acts as the classification branch to produce the classification results, i.e. $\mathtt{object}$ vs. $\mathtt{no\,object}\,(\emptyset)$. The $\emptyset$ class is used to represent that no object is detected, playing a similar role to the “background” class in the standard object detection approaches.
@@
Figure 2: Overview and detailed architecture design of our method. (Left) We utilize a DETR-inspired framework to predict the number of repetitive action cycles in an RGB video sequence. The architecture consists of a backbone network $\Phi(\cdot)$ for feature extraction, an encoder-decoder Transformer for processing query features, and prediction heads for action classification. By counting the queries classified as “repetitive actions”, we get the total count $N$. (Right) The detailed workflow within the black box on the left. Given the video features $\mathbf{F}$ as input, the encoder $\mathcal{E}(\cdot)$ produces two sets of query features, i.e. action queries and position queries. A query selection module then screens these queries, retaining only the most relevant ones for forwarding to the decoder $\mathcal{D}(\cdot)$. This process utilizes action classification results from prediction heads positioned after the encoder. The decoder adopts the Dynamic Action Query (DAQ) strategy, utilizing the selected action queries as priors, while incorporating a learnable position query and attending to the encoder’s output. The output embeddings from the decoder are then passed to their corresponding prediction head, which estimates the class label $\mathbf{A}$ and temporal location $\mathbf{P}$. During training, we employ bipartite matching to uniquely pair each prediction with a GT action cycle $\hat{\mathbf{P}}$, considering both class labels and temporal locations. Predictions that fail to match should yield $\emptyset$ class, signifying that they do not represent repetitive actions. The DAQ strategy addresses the open-set problem by contextually defining “repetitive actions” based on video features. To further distinguish these actions from $\emptyset$ distractions, we propose Inter-query Contrastive Learning (ICL). By integrating DAQ and ICL, our method effectively identifies contextually similar actions and excludes distractions, making it well-suited for tackling the challenges of the TRC problem.
@@ 3.2 Overview
Given an RGB video sequence with $T$ frames, the TRC task aims to predict an integer $N$ indicating the number of detected repetitive action cycles. Drawing inspiration from DETR [3], we streamline the problem as a set detection task in the temporal domain. The overall framework consists of three main components: a backbone network $\Phi(\cdot)$, an encoder-decoder Transformer, and prediction heads, as shown in Fig. 2. The backbone extracts video features $\mathbf{F}$, which is transformed to features of a set of action queries by an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$. Two prediction heads are added on top of the query features. The prediction heads consist of an action head and a position head. The action head classifies the query, while the position head estimates its temporal location.
Building on previous enhancements in DETR models [23], we introduce two distinct types of queries to enhance the query-to-feature similarity and accelerate convergence: action queries and position queries. Action queries are designed to capture features pertinent to action classification, whereas position queries concentrate on the temporal dimensions of an action. As depicted in the detailed diagram on the right side of Fig. 2, two specialized prediction heads (i.e. action head and position head) are tasked with processing these respective sets of queries to generate predictions. To further accelerate convergence, we introduce a post-encoder query selection module that acts as an initial sieve, filtering out the majority of queries unlikely to represent repetitive actions before forwarding the most promising ones to the decoder. This enhances the efficiency and focus of the subsequent prediction process.
@@
Considering the two challenges of the TRC tasks discussed in Sec. 1, we propose Dynamic Action Query (DAQ) strategy to address the open-set problem. DAQ directly uses dynamically updated action query features from the encoder to initialize the action query of the decoder. This strategy defines what “repetitive actions” should be detected based on the video content features in a dynamic and contextually aware manner. To tackle the second challenge, we define two action classes: “repetitive actions” and the “others ($\emptyset$)” class. We further propose Inter-query Contrastive Learning (ICL) to cluster repetitive action queries into a positive action set ($S^{+}$) and group the other queries to a negative set ($S^{-}$) in the feature space. Integrating DAQ and ICL allows our method to identify contextually similar action instances that are adaptively based on the video content, and exclude other distracting actions at the same time.
@@
Similar to DETR [3], during training, we employ bipartite matching which uniquely assigns a prediction to a GT action cycle. The matching procedure takes into account both the class prediction and the similarity of predicted and GT temporal locations. We then optimize the action cycle-specific losses. In inference, our method produces a fixed-size set of $Q$ predictions in a single pass, where $Q$ is set to be significantly larger than the typical number of action cycles in a $T$-frame video sequence. By counting the queries classified as “repetitive actions”, we get the final total count value $N$.
@@
In the following, we will introduce the network architecture design in Sec. 3.3, the DAQ and ICL modules in Sec. 3.4 and Sec. 3.5, and the model training in Sec. 3.6.
@@ 3.3 Model Architecture
Encoder. The encoder $\mathcal{E}(\cdot)$ is a classical Transformer [40] architecture which has ${L}^{\text{enc}}$ standard encoder layers. The encoder transforms the video features $\mathbf{F}$ into two distinct query types: action queries $\tilde{\mathbf{E}}^{\text{act}}\in\mathbb{R}^{T\times C}$ and position queries $\tilde{\mathbf{E}}^{\text{pos}}\in\mathbb{R}^{T\times C}$, as shown in Fig. 2 (right). This design enhances the query-to-feature similarity and accelerates convergence, where action queries capture features pertinent to action classification. On the other hand, position queries concentrate on the temporal dimensions of an action. We defer to the supplementary material the detailed encoder architecture.
@@
Query selection. We add a query selection module before passing the encoder output query features to the decoder. We route the query features outputted by the encoder to two respective prediction heads (described later in this section), which independently decode them into predictions for action class and temporal location, as depicted on the right side of Fig. 2. The action head processes the action queries $\tilde{\mathbf{E}}^{\text{act}}$ to estimate the action class $\tilde{\mathbf{A}}$ for each query, while the position head decodes the position queries $\tilde{\mathbf{E}}^{\text{pos}}$ to their temporal locations $\tilde{\mathbf{P}}$. Based on the prediction results, we discard query features classified as “others ($\emptyset$)”, retaining only those identified as “repetitive actions”. For the remaining queries, we rank them based on their classification confidence and preserve only the top $Q$ high-confidence queries. The others are discarded. After this selection process, both the action query features $\mathbf{E}^{\text{act}}$ and position query features $\mathbf{E}^{\text{pos}}$ have a dimension of $Q\times C$. Additional details will be provided when we discuss the prediction head later in this section.
@@
Decoder. The decoder $\mathcal{D}(\cdot)$ is also a classical Transformer [40] architecture, consisting of ${L}^{\text{dec}}$ standard decoder layers. It processes a set of action queries $\mathbf{Q}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, while simultaneously attending to the direct output (i.e. $\tilde{\mathbf{E}}^{\text{act}}$ and $\tilde{\mathbf{E}}^{\text{pos}}$) from the encoder, as depicted in Fig. 2 (right). Following the practice in DETR [3], these inputs are decoded in parallel across each decoder layer and transformed into the corresponding query features $\mathbf{D}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and $\mathbf{D}^{\text{pos}}\in\mathbb{R}^{Q\times C}$. The input position queries $\mathbf{Q}^{\text{pos}}$ are initialized as learnable parameters. To address the open-set challenge in the TRC task, we introduce the DAQ strategy, which initializes the decoder input action queries using the selected encoder embeddings $\mathbf{E}^{\text{act}}$, i.e. $\mathbf{Q}^{\text{act}}=\mathbf{E}^{\text{act}}$. The details of the DAQ strategy will be elaborated in Sec. 3.4. Using self- and encoder-decoder attention over these embeddings, our method globally reasons selected video frame features together while being able to use the whole frame-wise video feature as context. The two sets of $Q$ queries are then independently transformed into corresponding prediction results, i.e. an action head decodes $\mathbf{D}^{\text{act}}$ into the action class prediction $\mathbf{A}$, and a position head decodes $\mathbf{D}^{\text{pos}}$ into the temporal location prediction $\mathbf{P}$. We detail the two heads in the following subsection.
@@
Prediction heads. The prediction heads are designed to make the final predictions using the features processed by the encoder-decoder. They comprise two networks: an action head and a position head, both of which are Multi-Layer Perceptrons (MLPs).
@@
Bipartite matching. Following the practice of DETR [3], our method produces a fixed-size set of $Q$ predictions in a single pass, where $Q$ is set to be much larger than the typical number of action cycles in a $T$-frame video sequence. Note that the $Q$ predictions are in a random order. One of the main difficulties of training is to match the predicted action cycles (classes and temporal locations) to the GT action cycles, therefore, optimizing action cycle-specific losses, e.g. classification loss, temporal position loss, etc.
Let us denote GT repetitive action cycle labels by $\hat{\mathbf{Y}}=(\hat{\mathbf{A}},\hat{\mathbf{P}})$. $\hat{\mathbf{A}}=\mathbf{1}\in\mathbb{R}^{\hat{N}\times 1}$ denotes the $\hat{N}$ repetitive action class labels, where we assign the value 1 to the “repetitive actions”. $\hat{\mathbf{P}}\in\mathbb{R}^{\hat{N}\times 2}$ denotes the GT temporal positions of the $\hat{N}$ repetitive action, represented by $\hat{N}$ sets of midpoints and duration. Similarly, we denote the predicted repetitive actions by $\mathbf{Y}=(\mathbf{A},\mathbf{P})$, a set of size $Q$ predictions. We set $Q$ much larger than $\hat{N}$ empirically. We consider $\hat{\mathbf{Y}}$ also a set of size $Q$ padded with “others ($\emptyset$)”, i.e. we assign the value 0 to denote $\emptyset$ class. To find a bipartite matching between these two sets $\hat{\mathbf{Y}}$ and $\mathbf{Y}$, we search for a permutation of $Q$ elements $\sigma\in\mathfrak{S}_{Q}$ with the lowest cost:

### v4

@@ Abstract
Most existing methods rely on the similarity correlation matrix to characterize the repetitiveness of actions, but their scalability is hindered due to the quadratic computational complexity. In this work, we introduce a novel approach that employs an action query representation to localize class-agnostic repeated action cycles with linear computational complexity. Based on this representation, we develop two key components to tackle the essential challenges of temporal repetition counting. Firstly, to tackle open-set action counting, we define two action classes: “repetitive actions” and “others”. Instead of manually defining the repetitive action class, we propose a dynamic action query strategy. Here, each action query directly represents an extracted video feature, allowing the repetitive actions of interest to be dynamically defined based on the video content itself. Secondly, to distinguish these repetitive action queries from others, we propose inter-query contrastive learning. This performs contrastive clustering over the queries, pulling similar action patterns together while pushing apart those related to background or unrelated movements. As a result, queries classified as “repetitive actions” are considered as repetitive cycles, which are then used for counting. Thanks to the query-based representation and contrastive learning strategy, our method significantly outperforms previous works on accuracy while being more lightweight and time-efficient. On the challenging RepCountA benchmark, we outperform the state-of-the-art method TransRAC by 26.5% in OBO accuracy, with a 22.7% mean error decrease and 94.1% computational burden reduction. Code and models are publicly available at project page.
@@ 1 Introduction
Temporal periodicity is a ubiquitous phenomenon in the natural world. Temporal Repetition Counting (TRC) aims to accurately measure the number of repetitive action cycles within a given video and holds significant potential for applications such as fitness monitoring [fieraru2021aifit] and motion generation [zhu2023human].
@@
Pioneer methods  [laptev2005periodic, azy2008segmentation, cutler2000robust, tsai1994cyclic, pogalin2008visual, thangali2005periodic, chetverikov2006motion] represent time-series video data as one-dimensional signals and employ spectral analysis techniques such as the Fourier transform. While suitable for short videos with fixed periodic cycle lengths, these methods struggle to handle real-world scenarios with varying cycle lengths and sudden interruptions. Recent studies shift to deep learning-based methods [levy2015live, dwibedi2020counting, hu2022transrac, zhang2020context, li2024repetitive] and show promising performances. Notably, most of these methods, such as RepNet [dwibedi2020counting] and TransRAC [hu2022transrac], utilize a temporal similarity correlation matrix to depict repetitiveness, as illustrated in Fig.˜1 (a). Nevertheless, the computational complexity of this representation grows quadratically with the number of input frames $T$, highlighting a significant gap in scalability that hinders their application to real-world scenarios of varying action periods and dynamics.
Recent progress in action detection [shi2022react, liu2022end] introduces an efficient representation of action periods by associating each action instance with an action query, similar to DETR [carion2020end]. Inspired by this, we propose to formulate the TRC problem as a set prediction task where the goal is to detect every action cycle by representing it as an action query, as illustrated in Fig.˜1 (b). Based on this query representation, we use a Transformer encoder-decoder network [zhu2020deformable] to detect repetitive action instances and their temporal positions, defined by their midpoints and durations. This novel formulation reduces the complexity from quadratic to linear  [^1] In implementation, we employ deformable attention modules as proposed in DeformableDETR [zhu2020deformable]. and enables counting long videos with varying action periods. However, directly applying the action detection approaches [liu2022end, zhang2022actionformer, shi2022react] to the TRC problem proves inadequate (Tab.˜1) in addressing two distinctive challenges unique to TRC. These challenges underscore the complexity of TRC, highlighting why TRC is not merely another action detection task but requires a nuanced approach that considers the unique nature of repetitive actions. We highlight the two inherent differences between TRC and the classical action detection task:
@@
As a result, approaching TRC as a simple action detection task results in inferior performance as shown in Sec.˜4.3. In contrast, we propose two novel strategies to address these challenges and redefine the framework for TRC. In response to the first open-set challenge, we define two action classes: “repetitive actions” and “others ($\emptyset$)”. Instead of manually defining the repetitive action class, we propose a Dynamic Action Query (DAQ) strategy, which adaptively updates the action query using content features extracted from the video (Fig.˜2) This mechanism allows the decoder to attend to the repetitive actions based on the input video contents in a dynamic, contextually aware and class-agnostic manner. To tackle the second challenge, we further propose Inter-query Contrastive Learning (ICL), which ensures that primary repetitive action cycles of interest are grouped together in the learned representation space while being separated from other distractors, such as background noise. The integration of two core components (DAQ and ICL) ensures that the action instances are identified adaptively based on the video content and their contextual similarity, making our approach class-agnostic. In other words, our queries are designed to localize the contextually similar action instances, which aligns exactly with the definition of repetition counting. Extensive experiments validate the effectiveness of the two proposed designs.
@@
    We provide a novel perspective to tackle the TRC problem using a simple yet effective representation for action cycles, i.e. action query. Our approach reduces the computational complexity from quadratic to linear and is more lightweight and time-efficient.
@@
    Our method notably surpasses state-of-the-art (SOTA) methods in terms of both accuracy and efficiency on two challenging benchmarks. Notably, our method strikes an effective balance in handling various action periods and video lengths, offering a significant leap forward in the practical application of TRC technologies.
@@ 2.1 Temporal Repetition Counting
Traditional methods [laptev2005periodic, azy2008segmentation, cutler2000robust, tsai1994cyclic, pogalin2008visual, thangali2005periodic, chetverikov2006motion] frequently employ spectral or frequency domain techniques for the analysis of repetitive sequences, thereby preserving the underlying repetitive motion structures. While these conventional approaches are capable of effectively handling simple motion sequences or those characterized by fixed periodicity, they prove inadequate when confronted with non-stationary motion sequences encountered in real-world scenarios. In contrast, deep-learning-based approaches [levy2015live, dwibedi2020counting, hu2022transrac, zhang2020context, li2024repetitive] have demonstrated remarkable performance improvements. Notably, RepNet [dwibedi2020counting] and TransRAC [hu2022transrac] leverage temporal similarity matrices of actions to construct models for counting temporal repetitions. However, these similarity-matrix-based methods are not scalable for long videos due to their quadratic computational complexity. Another research line involves predicting the start and end points of each cycle [zhang2020context] from coarse to fine. Nevertheless, its practicality is hindered by the requirement for over 30 forward passes to count iteratively from a single video. In this paper, we introduce an effective action cycle representation by leveraging a Transformer encoder-decoder, which reduces the computational complexity from quadratic to linear and demonstrates superior performance in handling both fast and slow actions.
@@ 2.2 Temporal Action Detection
The field of temporal action detection [redmon2016you, zhao2017temporal, chao2018rethinking, zhang2022actionformer, lin2019bmn] is typically classified into two categories: anchor-based methods, and anchor-free methods. Anchor-based methods [zeng2019graph, li2021three, qing2021temporal] generate multiple anchors, subsequently classifying these anchors to determine the action boundaries. Anchor-free methods  [buch2019end, shou2017cdc, yuan2017temporal, lin2021learning] predict action instances by directly regressing the boundary and the center point of an action instance. With the rapid development of Transformer technology, DETR [carion2020end] is introduced for object detection task [zhu2020deformable, meng2021conditional, liu2022dab, zhang2022dino] and gains increasing popularity with promising performance. This paradigm promotes the study in many fields such as the action detection tasks [liu2022end, tan2021relaxed, vaswani2017attention, wang2021oadtr]. These methods establish a direct connection between action queries and the predicted action instances, enabling them to accurately predict the temporal boundaries of actions. Inspired by these promising results, we explore the possibility of utilizing a novel action query to represent the action cycle in TRC task. In contrast to existing action detection methods, our approach allows the model to capture the inherent repetitive content of an action cycle without relying on predefined class labels and effectively addresses confounding factors such as non-repetitive video backgrounds. This makes our approach well-suited for tackling the challenges of the TRC problem.
@@ 3.1 Preliminary
DETR [carion2020end] is a pioneering object detection framework that builds upon the Transformer encoder-decoder architecture [vaswani2017attention]. The overall DETR architecture [carion2020end] consists of three main components: a backbone to extract image features, an encoder-decoder Transformer, and the detection heads, i.e. feed-forward network (FFN) that makes the final detection prediction. The main features of DETR are the conjunction of a bipartite matching loss and transformers with (non-autoregressive) parallel decoding. The bipartite matching loss is a set-based Hungarian loss during training, which uniquely assigns a prediction to a GT object, and is invariant to a permutation of predicted objects. This design enables DETR to perform parallel processing and predict all objects simultaneously. We briefly review the workflow as follows.
Given the input image feature maps extracted by a CNN backbone, e.g. ResNet [he2016deep], DETR exploits a standard Transformer encoder-decoder network to transform the feature maps to be features of a set of object queries. An FFN and a linear projection are added on top of the object query features as the detection heads. The FFN acts as the regression branch that predicts the bounding box coordinates, i.e. box center coordinates, box height and width. The linear projection acts as the classification branch to produce the classification results, i.e. $\mathtt{object}$ vs. $\mathtt{no\,object}\,(\emptyset)$. The $\emptyset$ class is used to represent that no object is detected, playing a similar role to the “background” class in the standard object detection approaches.
@@
Figure 2: Overview and detailed architecture design of our method. (Left) We utilize a DETR-inspired framework to predict the number of repetitive action cycles given a video input. The framework consists of a backbone network $\Phi(\cdot)$ for feature extraction and the counting module (black box). The counting module, including an encoder-decoder Transformer for processing query features and prediction heads for action classification, outputs queries classified as “repetitive actions”, which are then counted to yield the total count NN. (Right) The detailed workflow of the counting module. Given the video features $\mathbf{F}$ as input, the encoder $\mathcal{E}(\cdot)$ produces a set of query features. A query selection module then screens these, retaining only the most relevant ones for forwarding to the decoder $\mathcal{D}(\cdot)$. The decoder $\mathcal{D}(\cdot)$ adopts the Dynamic Action Query (DAQ) strategy, using the selected action queries for initialization to dynamically define the “repetitive actions” class. The output embeddings from the decoder are then passed to their corresponding prediction head, which estimates the class label $\mathbf{A}$ and temporal location $\mathbf{P}$. During training, we employ bipartite matching to uniquely pair each prediction with a GT action cycle $\hat{\mathbf{P}}$. Predictions that fail to match are assigned the $\emptyset$ class. To further distinguish the “repetitive actions” from $\emptyset$ actions, we propose Inter-query Contrastive Learning (ICL). Finally, the queries classified as repetitive actions contribute to the total count NN.
@@ 3.2 Overview
Given an RGB video sequence with $T$ frames, the TRC task aims to predict an integer $N$ indicating the number of detected primary repetitive action cycles, whose class is not predefined. Drawing inspiration from DETR [carion2020end], we streamline the problem as an open-set detection task in the temporal domain and propose to use the action query to represent each potential repetitive cycle. The overall framework consists of two main components: a backbone network $\Phi(\cdot)$ and the counting module. The backbone extracts video features $\mathbf{F}$. The counting module, depicted as the black box in Fig.˜2, is composed of a Transformer encoder $\mathcal{E}(\cdot)$, a decoder $\mathcal{D}(\cdot)$, and prediction heads. The extracted $\mathbf{F}$ are transformed into a set of action queries by the Transformer encoder and decoder. Two prediction heads are then added on top of these query features, aiming to classify each query and estimate its temporal location. Once an action query is classified as a repetitive cycle, the action count $N$ increases by 1.
@@
Considering the two challenges of the TRC tasks discussed in Sec.˜1, we propose Dynamic Action Query (DAQ) strategy to address the open-set problem. Specifically, we define two action classes: “repetitive actions” and the “others ($\emptyset$)” class. Instead of manually defining the specific class label for the “repetitive actions”, DAQ directly leverages dynamically updated video features from the encoder to initialize the action query for the decoder. This enables the adaptive definition of “repetitive actions” based on the video content, in a contextually aware and class-agnostic manner. To tackle the second challenge of distinguishing identical repetitive actions, we further propose Inter-query Contrastive Learning (ICL). This approach clusters queries representing identical repetitive actions into a positive action set ($S^{+}$) and groups the other queries into a negative set ($S^{-}$) in the feature space. Integrating DAQ and ICL allows our method to identify class-agnostic similar action instances that are adaptively based on the video content, and exclude other distracting actions at the same time.
@@
Similar to DETR [carion2020end], during training, we employ bipartite matching which uniquely assigns a prediction to a GT action cycle. The matching procedure takes into account both the class prediction and the similarity of predicted and GT temporal locations. We then optimize the action cycle-specific losses. In inference, our method produces a fixed-size set of $Q$ predictions in a single pass, where $Q$ is set to be significantly larger than the typical number of action cycles in a $T$-frame video sequence. By counting the queries classified as “repetitive actions”, we get the final total count value $N$.
@@
In the following, we will introduce the network architecture design in Sec.˜3.3, the DAQ and ICL modules in Sec.˜3.4 and Sec.˜3.5, and the model training in Sec.˜3.6.
@@ 3.3 Model Architecture
As illustrated in Fig.˜2, our model can be divided into two main components: a backbone $\Phi(\cdot)$ and the counting module (black box). The backbone extracts features $\mathbf{F}$ from the raw video input. The counting module then takes these video features $\mathbf{F}$ as input and ultimately outputs the count N. Specifically, the counting module includes a Transformer encoder-decoder, a query selection module, prediction heads, and a bipartite matching module. Next, we introduce each of them in detail.
@@
Encoder. The encoder $\mathcal{E}(\cdot)$ is a classical Transformer [vaswani2017attention] architecture which has ${L}^{\text{enc}}$ standard encoder layers. The encoder transforms the video features $\mathbf{F}$ into two distinct query types: action queries $\tilde{\mathbf{E}}^{\text{act}}\in\mathbb{R}^{T\times C}$ and position queries $\tilde{\mathbf{E}}^{\text{pos}}\in\mathbb{R}^{T\times C}$, as shown in Fig.˜2 (right). Action queries capture features pertinent to action classification, while position queries concentrate on the temporal dimensions of an action. These features prepare the decoder for action query initialization, from which it will identify repetitive action cycles. Please refer to the supplementary material for the detailed encoder architecture.
@@
Query selection. We add a query selection module before passing the encoder output query features to the decoder, as depicted on the right side of Fig.˜2. This module aims to perform an initial filtering over the encoded video features $\mathbf{F}$. Since the encoder $\mathcal{E}(\cdot)$ produces a large number of tokens $\tilde{\mathbf{E}}^{\text{act}}\in\mathbb{R}^{T\times C}$ and $\tilde{\mathbf{E}}^{\text{pos}}\in\mathbb{R}^{T\times C}$, many of them may correspond to background or distracting content. Therefore the query selection module selects a subset of $Q$ informative features from the original $T$ tokens as $\mathbf{E}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and $\mathbf{E}^{\text{pos}}\in\mathbb{R}^{Q\times C}$. In practice, we route the query features outputted by the encoder to two respective prediction heads (described later in this section), which independently decode them into predictions for action class and temporal location. The action head processes the action queries $\tilde{\mathbf{E}}^{\text{act}}$ to estimate the action class $\tilde{\mathbf{A}}$ for each query, while the position head decodes the position queries $\tilde{\mathbf{E}}^{\text{pos}}$ to their temporal locations $\tilde{\mathbf{P}}$. Based on the prediction results, we discard query features classified as “others ($\emptyset$)”, retaining only those identified as “repetitive actions”. For the remaining queries, we rank them based on their classification confidence and preserve only the top $Q$ high-confidence queries. The others are discarded. After this selection process, both the action query features $\mathbf{E}^{\text{act}}$ and position query features $\mathbf{E}^{\text{pos}}$ have a dimension of $Q\times C$. Additional details will be provided when we discuss the prediction head later in this section.
@@
Decoder. The decoder $\mathcal{D}(\cdot)$ is also a classical Transformer [vaswani2017attention] architecture, consisting of ${L}^{\text{dec}}$ standard decoder layers. It processes a set of action queries $\mathbf{Q}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, while simultaneously attending to the direct output (i.e. $\tilde{\mathbf{E}}^{\text{act}}$ and $\tilde{\mathbf{E}}^{\text{pos}}$) from the encoder, as depicted in Fig.˜2 (right). Following the practice in DETR [carion2020end], these inputs are decoded in parallel across each decoder layer and transformed into the corresponding query features $\mathbf{D}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and $\mathbf{D}^{\text{pos}}\in\mathbb{R}^{Q\times C}$. The input position queries $\mathbf{Q}^{\text{pos}}$ are initialized as learnable parameters. To address the open-set challenge in the TRC task, we introduce the DAQ strategy, which initializes the decoder input action queries using the selected encoder embeddings $\mathbf{E}^{\text{act}}$, i.e. $\mathbf{Q}^{\text{act}}=\mathbf{E}^{\text{act}}$. The details of the DAQ strategy will be elaborated in Sec.˜3.4. Using self- and encoder-decoder attention over these embeddings, our method globally reasons selected video frame features together while being able to use the whole frame-wise video feature as context. The two sets of $Q$ queries are then independently transformed into corresponding prediction results, i.e. an action head decodes $\mathbf{D}^{\text{act}}$ into the action class prediction $\mathbf{A}$, and a position head decodes $\mathbf{D}^{\text{pos}}$ into the temporal location prediction $\mathbf{P}$. We detail the two heads in the following subsection.
@@
Prediction heads. The prediction heads are designed to make the final predictions using the features processed by the encoder-decoder, as shown in Fig.˜2 (right). They comprise two networks: an action head and a position head, both of which are Multi-Layer Perceptrons (MLPs).
@@
Bipartite matching. Note that our method produces a fixed-size set of $Q$ predictions in a random order, where $Q$ is set to be much larger than the typical number of action cycles in a $T$-frame video sequence. To enforce supervision, we employ a matching strategy to pair the predicted action cycles (classes and temporal locations) with the GT action cycles, as illustrated in the bottom right of Fig.˜2.
[…]